A mutual fund manager has a $20 million portfolio with a beta of 1.80. The risk-free rate is 7.75%, and the market risk premium is 6.5%. The manager expects to receive an additional $5 million, which she plans to invest in a number of stocks. After investing the additional funds, she wants the fund's required return to be 16%. What should be the average beta of the new stocks added to the portfolio? Negative value, if any, should be indicated by a minus sign. Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1
As per CAPM
expected return = risk-free rate + beta * (Market risk premium)
Expected return% = 7.75 + 1.8 * (6.5)
Expected return% = 19.45
Total New portfolio value = Value of Old portfolio + Value of Add. Inv
=20+5
=25
Weight of Old portfolio = Value of Old portfolio/Total New portfolio Value
= 20/25
=0.8
Weight of Add. Inv = Value of Add. Inv/Total New portfolio Value
= 5/25
=0.2
Exp return of New portfolio = Weight of Old portfolio*Exp return of Old portfolio+Weight of Add. Inv*Exp return of Add. Inv
16 = 19.45*0.8+Exp return of Add. Inv*0.2
Exp return of Add. Inv = 2.2
As per CAPM
expected return = risk-free rate + beta * (Market risk premium)
2.2 = 7.75 + Beta * (6.5)
Beta = -0.85